OpenStudy (anonymous):

The line with m = 3 and b = -2

OpenStudy (anonymous):

y=3x-2 if you want it in y=mx+b form.

OpenStudy (amoodarya):

\[y=mx+b \] m= slope b is y-intercept like this y-intercept =6 slope is -3 so y=mx+b so y=-3x+6 now do like this

OpenStudy (anonymous):

thanks